# Measurement-induced phase transition in periodically driven free-fermionic systems

### v1

@@ Appendix A $T=100$ result for square pulse protocol
Here, we discuss the $T=100$ results. Dashed lines correspond to no-drive results for $J=1/2$. It shows that almost up to $L\simeq 100$, the no-drive data is the same as the $T=100$ data. Afterward, $T=100$ data starts bending, while no-drive data keeps on increasing. In the main text, we show that in the case of square pulse protocol and a given driving time period $T$, for $L>>L_{c}$, the steady state entanglement entropy does not change with $L$. The value of $L_{c}$ was estimated using fitting function  (16). It turns out that $L_{c}$ increases exponentially with $T$, the fitted function predict the $L_{c}$ to be ${}10^{16}$ for $T=100$ and $\gamma=0.05$. This suggests that due to our limitations of the accessible system size, we will be unable to identify the area-law phase for $\gamma=0.05$ and $T=100$. As expected, Fig. 12 shows no trace of area law for $\gamma=0.05$. The data can mislead us to predict the usual BKT type critical to area law entanglement phase transition as we increase $\gamma$.
@@
Appendix B Cost-function minimization technique
In order to see the BKT transition, we had to perform a finite-size scaling analysis. The job is to find out the $\gamma_{c}$ for which one can get the best possible data-collapse as observed in Fig. 7 (lower panel). We take $X_{i}=(\gamma-\gamma_{c})(\ln L)^{2}$ for all the values of $\gamma$ and $L$, and $Y_{i}=\overline{S(L,\gamma)}-\overline{S(L,\gamma_{c})}$ sorted according to the increasing values of $X_{i}$. The values of $\overline{S(L,\gamma)}$, which are not in the list, have been calculated using linear extrapolation. Now, one can define the cost function as Šuntajs et al. (2020),
@@ Appendix C F-test
where $\tilde{E}_{L}$ ($\tilde{E}_{\ln(L)}$) is the sum of the squared errors between $\overline{S(L)}$ and $\tilde{S}_{L}$ ($\tilde{S}_{\ln(L)}$). The way $F_{T}$ is constructed, a lower $F_{T}$ will favor volume law, and a higher $F_{T}$ will favor logarithmic scaling. We exercise this test for our $\overline{S(L)}$ vs $L$ data and found the $F_{T}>>1$ for $T=5$, $20$, $100$, $150$, $200$ and for $\gamma=0.05$, and $\epsilon=3J/4$. It confirms that the scaling is most likely logarithmic.

### v2

@@ Appendix A $T=100$ result for square pulse protocol
Here, we discuss the $T=100$ results. In Fig. 18, the dashed lines correspond to no-drive results for $J=1/2$. It shows that almost up to $L\simeq 100$, the no-drive data is the same as the $T=100$ data. Afterward, $T=100$ data starts bending, while no-drive data keeps on increasing. In the main text, we show that in the case of square pulse protocol and a given driving time period $T$, for $L>>L_{c}$, the steady state entanglement entropy does not change with $L$. The value of $L_{c}$ was estimated using fitting function  (20). It turns out that $L_{c}$ increases exponentially with $T$, the fitted function predict the $L_{c}$ to be ${}10^{16}$ for $T=100$ and $\gamma=0.05$. This suggests that due to our limitations of the accessible system size, we will be unable to identify the area-law phase for $\gamma=0.05$ and $T=100$. As expected, Fig. 18 shows no trace of area law for $\gamma=0.05$. The data can mislead us to predict the usual BKT type critical to area law entanglement phase transition as we increase $\gamma$.
Appendix B Cost-function minimization technique
@@
Figure 20: Cost function $F_{Y}$ vs. $\gamma_{c}$ plot following the scaling Eq.  18 for $\epsilon=3J/4,T=5$ and no drive $J=1/2$. The red star represents the optimal $\gamma_{c}$ for $\epsilon=3J/4$, and the black circle represents the optimal $\gamma_{c}$ value for $J=1/2$ no drive.
In order to see the BKT transition, we had to perform a finite-size scaling analysis. The job is to find out the $\gamma_{c}$ for which one can get the best possible data-collapse as observed in Fig. 7 (lower panel). We take $X_{i}=(\gamma-\gamma_{c})(\ln L)^{2}$ for all the values of $\gamma$ and $L$, and $Y_{i}=\overline{S(L,\gamma)}-\overline{S(L,\gamma_{c})}$ sorted according to the increasing values of $X_{i}$. The values of $\overline{S(L,\gamma)}$, which are not in the list, have been calculated using linear interpolationFuji and Ashida (2020); Skinner et al. (2019b). Now, one can define the cost function as Šuntajs et al. (2020b),
@@ Appendix C F-test
where $\tilde{E}_{L}$ ($\tilde{E}_{\ln(L)}$) is the sum of the squared errors between $\overline{S(L)}$ and $\tilde{S}_{L}$ ($\tilde{S}_{\ln(L)}$). The way $F_{T}$ is constructed, a lower $F_{T}$ will favor volume law, and a higher $F_{T}$ will favor logarithmic scaling. We exercise this test for our $\overline{S(L)}$ vs $L$ data and found the $F_{T}>>1$ for $T=5$, $20$, $100$, $150$, $200$ and for $\gamma=0.05$, and $\epsilon=3J/4$. It confirms that the scaling is most likely logarithmic(for finite size).
Appendix D Effective central charge in the absence of drive
In this section, we present the results for the effective central charge in the absence of driving with $J=1/2$. Though these are the no-drive results, for the accessible system sizes in our study, the high-frequency limit ($T<1$) results should resemble them. This is because the entanglement entropy data in that limit is almost indistinguishable from the no-drive data with $J_{eff}=\epsilon/2$. Using the same cost function minimization technique, we get the transition point $\gamma_{c}=0.148\pm 0.039$. We calculate the $l^{*}_{c}$ for which $c^{eff}$ is maximum for some $\gamma>\gamma_{c}$, and best fit gives $l^{*}_{c}\sim\gamma^{-2.01}$, from the fitting we calculate the $l^{*}_{c}$ values and try to match with the actual simulated data with system size $1004$ for $\gamma\leq\gamma_{c}$ as we did in the main text for $T=5$. In Fig. 21, we observe evidence of bending for $\ \gamma=0.12$ and $0.1$, but for $\gamma=0.07$, the bending is not very clearly evident. Note that in Ref. Starchl et al. (2024a), a similar analysis was done for a different measurement protocol.
Figure 21: Effective central charge $c^{eff}_{l}$ vs. $l$ plots for different $\gamma$ values for nodrive, $J=1/2$. Black circular points represent the maximum of the central charge and red star points represent the position of $l^{*}_{c}$ for which $c^{eff}_{l}$ is the maximum estimated from the fitting. Inset shows the best fit, which gives $l^{*}_{c}\sim\gamma^{-2.01}$. The maximum system size is taken as $L=1004$.
Appendix E QR decomposition and normalization in each trotter step
The problem discussed in the paper $U$ evolves as,
$U(t+dt)=e^{M}e^{-iH(t)dt}U(t),$
where $H(t)=J(t)\sum_{i=1}^{L}(\hat{c}_{i}^{{\dagger}}\hat{c}_{i+1}+h.c.)$ is the single-particle Hamiltonian. Under this non-hermitian evolution, the isometry $U(t+dt)^{\dagger}U(t+dt)=\mathcal{I}_{N\crossproduct N}$ does not satisfy. To recover this isometry property, we do $QR$ decomposition as, $U(t+dt)=QR$. Where $Q$ is an $L\crossproduct N$ matrix with orthonormal columns, satisfying
$Q^{\dagger}Q=\mathcal{I}_{N\times N}.$
and $R$ is an $N\times N$ upper triangular matrix. Since $Q$ retains the orthonormality condition, we replace $U(t+dt)$ with $Q$, ensuring:
$U(t+dt)\rightarrow Q,\quad\text{where }Q^{\dagger}Q=\mathcal{I}_{N\times N}.$
Thus, the QR decomposition projects $U(t+dt)$ back onto the manifold of isometric matrices, maintaining the normalization of the quantum state.
Figure 22: Fluctuation vs. $\gamma$ for different subsystem size $l$ for no drive(Upper panel) and $T=5$ case(lower panel).
Since the columns of $Q$ remain orthonormal, it ensures that the quantum state remains normalized.
Appendix F Fluctuation of steady state entanglement entropy
Figure 23: Upper left panel: Scaling of standard deviation/fluctuation of long time averaged entanglement entropy over various trajectories $\Delta S$ with subsystem size $l$ for different driving periods $T$ for $\gamma=0.05$. Upper right panel: fluctuation-mean ratio vs cord length for different driving periods. System size $L=600$ for $\gamma=0.05$. The lower panel presents the result for $\gamma=1$.
In this section, we have calculated the standard deviation of the time-averaged entanglement entropy for various different trajectories $\Delta S=\sqrt{{\overline{S^{2}}}-(\overline{S})^{2}}$, where the overline represents the trajectory average and $S$ is the long time averaged(taken stroboscopically) entanglement entropy value of each trajectory. . From Fig. 22 we see for small $l$ (sub-system size), the fluctuation $\Delta S$ tends to show a peak near the critical $\gamma_{c}$; however, for reasonably large $l$, that peak seems to disappear and the $\Delta S$ seems to decrease with $\gamma$ monotonically. This signature prevails for the no-drive and even for the finite $T=5$ case.
More precisely, we investigate how the standard deviationKalsi et al. (2022) of the long-time averaged entanglement entropy over various trajectories behaves with system sizes for different $T$, In Fig. 23 upper left panel, we see that in general, fluctuation($\Delta S$) increases with $l$, and also for small system size, it increases with $T$; for large $l$, for different $T$, the difference between their fluctuations is getting smaller for $\gamma=0.05$. Also, we calculated the fluctuation and mean ratio in the upper right panel for $\gamma=0.05$, which decreases with $l$, and for a given small sub-system size, decreases with decreasing $T$. Again, in the large $l$ limit, the difference between the fluctuation-mean ratio gets smaller for different $T$s. On the other hand, the lower panels show in the area-law regime $\Delta S$ and the fluctuation-mean ratio both are not scaling with $l$, but they increase with driving period $T$ for $\gamma=1$.